OpenStudy (anonymous):

What is the solution of √5x+1-√x=5? x = 0 x = 16 and x = 0 x = 16 x = 16 and x = 1

OpenStudy (anonymous):

is the equation this? \[\sqrt{5x} +1 -\sqrt{x} = 5\] or \[\sqrt{5}*x +1 -\sqrt{x} = 5\]

OpenStudy (raden):

16